OpenStudy (kearsten):

fan and medal number of solutions 0=x^2-5x

OpenStudy (kearsten):

@Hero @nincompoop @Preetha @iambatman @triciaal

OpenStudy (anonymous):

0 = x^2 - 5x 0 = x (x-5) x= 0 or x = 5